A 15-kilowatt motor is used to hoist an 800 kg bucket of concrete to the 20th floor of a building under construction to height of 90 m. assuming no energy is lost along the process, how much time is required?


1
Expert's answer
2021-01-27T10:20:28-0500

"work done, mgh=power (P)\times time(t)"

"mgh=P\times t"

"800\times9.8\times90=15\times1000 \times t"

"t=(800\times9.8\times90)/15000=47.04seconds"